OpenStudy (amriju):

ok..for this question you will need to tell me the process too: let x( x is a multidigit number) be a palindrome number. Find the least x such that x(x+1) is also a palindrome number. Any valid process will be given a medal...merely posting answers won't do

OpenStudy (amriju):

Process need not be elaborate..just the hint will do..or the algo

OpenStudy (anonymous):

since we have any 1digit number is a palindrome number, so that you can have x =1 and it satisfies 1(1+1) =2, right? That is the "least" x , right?

OpenStudy (amriju):

palindrome my friend...x has to be a multi digit number...you know what a palindrome number is?

OpenStudy (anonymous):

Actually, I don't know what it is but your problem is interesting so that I made a short research about it. And I know what it is , how it works http://en.wikipedia.org/wiki/Palindromic_number However, if x is multidigit number, let me work more. Hopefully, I can figure out something or at least I learn something new.

OpenStudy (amriju):

its quite a good question you know...damn interesting one...if you are interested in number theory

OpenStudy (anonymous):

if it is number theory, then @ganeshie8 is the best

OpenStudy (anonymous):

By testing, I got 77*(77+1) =6006 , that is the solution, but not know how to put it in logic yet. hihihi..However, we can derive the process from the result, right? (hopefully)

OpenStudy (amriju):

@ganeshie8 ...can u solve this?

OpenStudy (amriju):

@OOOPS ...the answer is easy..the process is what matters

OpenStudy (anonymous):

@amriju yes, you are right, let wait for him; he is asleep; just few more hours, he wakes up and help us. hihihi

OpenStudy (puzzler7):

A brute force method is a legit process to do this.

OpenStudy (amriju):

nope...thats the beauty..u dont need brute force..:)

OpenStudy (puzzler7):

But it is a legit method, just not the best.

OpenStudy (amriju):

solve it mathematically..not by those methods...i want that..brute force will be merely testing each value..

OpenStudy (anonymous):

If we restrict \(x\) to be a two-digit number, you can write \[x=10x_2+x_1\] where \(x_1\) and \(x_2\) are integers between 0 and 9 (including 0 and 9, except \(x_2\not=0\)). For \(x\) to be palindromic, you must have \(x_1=x_2\), so \(x=11x_2\). From here you can brute force it and determine that \(x_2=7\), or \(x=77\) is the first palindrome to yield a palindromic \(x(x+1)\), but we want a more elegant method, yes? I can't promise perfection, but I'll think about this problem...

OpenStudy (amriju):

even from there you dont need a brute force..

OpenStudy (amriju):

I'll post the answer if you want me to..one of the process that requires no prerequisites...

OpenStudy (anonymous):

I tried going about this by letting \(y=x(x+1)\), then either \[y=100y_3+10y_2+y_1~~~~\text{or}~~~~y=1000y_4+100y_3+10y_2+y_1\] Equating this with the \(x\) expression, \[(10x_2+x_1)(10x_2+x_1+1)=100y_3+10y_2+y_1\\~~~~~~~~~~~~~~~~~~~~~~~~\text{or}\\ (10x_2+x_1)(10x_2+x_1+1)=1000y_4+100y_3+10y_2+y_1\] depending on the value of \(x_1\). Then I tried solving for \(x_1=x_2\) by matching up the digits places. No luck though, unless there's something I'm not seeing.

ganeshie8 (ganeshie8):

for the last case, \(\large x_1 = x_2, y_4 = y_1, y_3=y_2\) give us a diaphontine equation : \[\large x_1(11x_1+1) = 91y_1+10y_2\] solving this also gives us cases

ganeshie8 (ganeshie8):

so the easy way is to bruteforce and test the 8 cases : 11,22,..,88